# NuHepMC: A standardized event record format for neutrino event generators Steven Gardiner1, Joshua Isaacson1 and Luke Pickering2

### v2

@@ 1 Introduction
Recent discussions about the future of neutrino research, both at workshops focused on simulations of neutrino interactions [3] and as part of the HEP-wide Snowmass 2021 community planning process [4], have emphasized the need for greater flexibility in the use of neutrino event generators and related software. A particularly problematic technical barrier is the current lack of a common data format for representing the output events: lists of simulated particles involved in a neutrino interaction together with information describing their properties and relationships. At present, each neutrino event generator group maintains a unique output format, substantially complicating the use of multiple generators in large-scale experimental simulation workflows [5, 6]. The adoption of standard formats in the collider community has streamlined many components of the analysis pipeline. This has enabled straightforward analysis preservation in tools like Rivet [7, 8], allowed theorists to reinterpret experimental limits on one exotic physics scenario in light of others [9, 10], simplified comparisons between generators (see, e.g., Ref. [11]), and supported interoperability between simulation tools [12, 13, 14].
@@
In Sec. 2, we present the specification of the NuHepMC format. The flux-averaged total cross section, a particularly important quantity for analyzing neutrino scattering simulations, is discussed in Sec. 3. Finally, Sec. 4 uses NuHepMC as an output format for GENIE, NuWro, NEUT, ACHILLES [23], and MARLEY [24] as well as an input format in NUISANCE to demonstrate a first NuHepMC-based analysis of neutrino event generator predictions.
@@ 2 Specification
The details of the NuHepMC specification are broken down into three categories that describe four components from HepMC3. Each element of the specification is labeled as a Requirement, a Convention, or a Suggestion. The requirements dictate a minimum level of information to be included when writing out events. The conventions are optional details that an event generator group can choose to include or omit while still conforming to the NuHepMC standard. Finally, the suggestions are optional recommendations that are less strongly encouraged than the conventions.
@@ 2.1 Labeling scheme
The elements of the specification are enumerated below in the form <Component>.<Category>.<Index>, where the component of interest is given as G for generator run metadata, E for event metadata, V for vertex information, and P for particle information. The category is denoted by R for a requirement, C for a convention, and S for a suggestion. For example, the second convention for event metadata would be labeled as E.C.2.
If conventions or suggestions prove useful, become widely adopted, and are considered stable, they may become requirements in future versions of this specification. G.C.1 defines a convention for event generator authors to signal whether or not specific optional elements of this specification have been followed.
@@ 2.2 HepMC3 C ++ classes
Throughout this standard, references are made to various HepMC3 C ++ classes, e.g., HepMC3::GenRunInfo. However, these are used as a convenient handle for data objects. This specification should not be considered specific to the HepMC3 C ++ reference implementation.
@@ 2.3 Generator Run Metadata
All NuHepMC Vectors must contain a HepMC3::GenRunInfo instance. -
@@
A NuHepMC HepMC3::GenRunInfo instance must contain the following attributes that specify the version of NuHepMC that is implemented:-
type: HepMC3::IntAttribute,
name: "NuHepMC.Version.Major" -
type: HepMC3::IntAttribute,
name: "NuHepMC.Version.Minor" -
type: HepMC3::IntAttribute,
name: "NuHepMC.Version.Patch"
This document describes version 0.9.0 of NuHepMC.
@@ A NuHepMC HepMC3::GenRunInfo instance must contain a
A NuHepMC HepMC3::GenRunInfo instance must contain a HepMC3::VectorIntAttribute named "NuHepMC.ProcessIDs" listing all physics process IDs as integers. For each valid process ID, the HepMC3::GenRunInfo instance must also contain two other attributes giving a name and description of each:-
@@
where ID enumerates all process IDs present in "NuHepMC.ProcessIDs". (See also E.C.1).
G.R.5
@@
HepMC3::VectorIntAttribute named "NuHepMC.VertexStatusIDs" declaring any generator-specific status codes used. Including the standard HepMC3 codes in this list is optional, but they must not be reused to mean something different than in the HepMC3 specification. For each declared vertex status, the HepMC3::GenRunInfo instance must also contain two other attributes giving a name and description of each:-
@@
where ID enumerates all status codes present in "NuHepMC.VertexStatusIDs" (See also V.R.1).
G.R.6
@@
HepMC3::VectorIntAttribute named "NuHepMC.ParticleStatusIDs" declaring any generator-specific status codes used. Including the standard HepMC3 codes in this list is optional, but they must not be reused to mean something different than in the HepMC3 specification. For each valid particle status, the HepMC3::GenRunInfo instance must also contain two other attributes giving a name and description of each:-
@@
where ID enumerates all status codes present in "NuHepMC.ParticleStatusIDs" (see P.R.1 for more details).
G.R.7
Event Weights:
For weights that will be calculated for every event, HepMC3 provides an interface for storing the weight names only once in the HepMC3::GenRunInfo instance. At least one event weight named "CV" must be declared in the HepMC3::GenRunInfo instance and filled for every event.This weight may be 1 or constant for every event in a generator run (in the case of an unweighted event vector). This weight must always be included by a user when producing correctly-normalized predictions from a NuHepMC Vector and must not be assumed to be always 1. The exact form of this weight and whether it is the only information required to properly normalize a prediction are considered implementation details.
G.R.8
@@
Essentially all event generators in HEP use a standard set of integer codes to identify particle species. This numbering scheme is maintained by the Particle Data Group (PDG) and is regularly updated in their Review of Particle Physics [25, Sec. 45, p. 733].We expect that neutrino event generators may need to use codes for non-standard particle species (i.e., those without an existing PDG code) for a variety of applications. This could include simulating exotic physics processes involving new particles as well as implementing bookkeeping methods involving generator-specific quasiparticles.
@@
HepMC3::VectorIntAttribute named "NuHepMC.AdditionalParticleNumbers" declaring any particle codes used that are not defined in the current PDG numbering scheme. Including any of the standard codes in this list is permitted but not required. The standard particle codes must not be reused to mean something different than in the PDG specification.For each additional particle code, the HepMC3::GenRunInfo instance must also contain an attribute giving a unique name to the represented particle species:
@@
"NuHepMC.AdditionalParticleNumbers".See also G.C.8 for a suggested way of storing descriptions of these special particle species.
@@
Signaling Followed Conventions:
To signal to a user that an implementation follows a named convention from this specification, a HepMC3::VectorStringAttribute should be added to the HepMC3::GenRunInfo instance named "NuHepMC.Conventions" containing the names of the conventions adhered to. -
@@
Vector Exposure (Standalone):
Each Vector should contain a description of the exposure of the generator run. When running a standalone event simulation this will often correspond to the number of events requested, which may differ from the number of events output in cases where events that are not written out must contribute to the total cross section calculation.-
type: HepMC3::LongAttribute,
name: "NuHepMC.Exposure.NEvents"
Implementations should not adhere to both G.C.2 and G.C.3 simultaneously.
@@
Vector Exposure (Experimental):
Each Vector should contain a description of the exposure of the generator run. When simulating with some experimental exposure, often represented for accelerator neutrino experiments in units of “protons on target” (POT), the exposure should be described. Two attributes are reserved for signaling the exposure used to users. One or both can be provided.-
type: HepMC3::DoubleAttribute, name: "NuHepMC.Exposure.POT"
type: HepMC3::DoubleAttribute, name: "NuHepMC.Exposure.Livetime"
Implementations should not adhere to both G.C.2 and G.C.3 simultaneously.
G.C.4
Cross Section Units and Target Scaling:
There are a variety of units typically used to report both measured and predicted cross sections in HEP. For neutrino cross sections specifically, per nucleon is common, but not ubiquitous. We want to provide a sensible recommended default while preserving the flexibility for an implementation to signal a different choice. One or both of the following HepMC3::StringAttributes may be included on the HepMC3::GenRunInfo to indicate the cross section units used within a vector.-
"NuHepMC.Units.CrossSection.Unit". Possible values of the attribute are not restricted, but we reserve the meanings of the following:
"pb": Picobarns or . This is our recommended default.
"1e-38 cm2": The choice of in this option is the most frequent in the neutrino literature.
"NuHepMC.Units.CrossSection.TargetScale". Possible values of the attribute are not restricted, but we reserve the meanings of the following for ease of compatibility with existing conventions:
"PerTargetAtom": Our recommendation. Choosing “atom” rather than “nucleus” in this context removes ambiguity when considering neutrino interactions with atomic electrons.
"PerTargetMolecule": Sometimes used for hydrocarbon- and water-target measurements.
"PerTargetNucleon": A common choice in the literature.
"PerTargetMolecularNucleon": Another common choice in the literature and in existing neutrino event generators. We recommend against implementations using this scheme.
If this convention is signalled, the chosen units should be assumed to apply to cross section information stored according to G.C.5, E.C.4, E.C.2, and E.C.3. For further discussion of the schemes detailed above and the motivation for recommending against "PerTargetMolecularNucleon", see A.
It is ultimately up to the user to parse these attributes and decide whether any additional scaling is needed for their purposes. If these attributes are not present, then the cross section should be assumed to be in picobarns per target atom. We strongly recommend that implementations use this default.
G.C.5
Flux-averaged Total Cross Section:
The flux-averaged total cross section, , is a scaling factor that is needed to convert a distribution of simulated events into a prediction of the flux-averaged cross-section—an experimentally-accessible quantity. Details on the definition of this quantity are given in Sec. 3.1.The value of is not always known at the beginning of a generator run. As described in B, a running estimate of the flux-averaged total cross section may be computed as events are generated. Element E.C.4 of this specification provides a means of storing the value of this running estimate in each event. If known at the start of a run, the value of should be stored as a HepMC3::DoubleAttribute in the HepMC3::GenRunInfo named "NuHepMC.FluxAveragedTotalCrossSection".
G.C.6
@@
Modelling components implemented based on published work should always be fully cited. The HepMC3::GenRunInfo should contain at least one
HepMC3::VectorStringAttribute for each relevant modelling component, named according to the pattern "NuHepMC.Citations.Comp.Type". Valid substitutions for the Comp and Type fields are not restricted by this standard beyond the requirement that they are pure mixed-case alpha-numeric. We suggest using Comp=Generator for specifying the main citation for the interaction generator and Comp=Process[ID] for individual processes. For common reference formats in the HEP field, we suggest some common values for the Type field:-
@@
We hope that automatic bibliography generation tools using this metadata will be built.
G.C.7
@@
Each vector should contain a description of the beam particle flux used to simulate the output event vector. For many truth studies and experimental simulations where the detector is not physically close to the source, a simple beam energy distribution is enough to describe the particle beam. The two types of energy distribution covered by this convention are mono-energetic beams and those with distributions described by a simple histogram. The type should be signalled via a HepMC3::StringAttribute named "NuHepMC.Beam.Type" with value "MonoEnergetic" or "Histogram" stored on the HepMC3::GenRunInfo. For both types, relevant units can be signalled via two attributes:-
"NuHepMC.Beam.EnergyUnit". Possible values of the attribute are not restricted, but we reserve the meanings of "MEV" and "GEV". This attribute should always exist and be not empty.
"NuHepMC.Beam.RateUnit". Possible values of the attribute are not restricted, but we reserve the meaning of "Arbitrary" to signal that the normalization of the distribution was not known or used by the simulation. If this attribute is not used then the normalization will be assumed arbitrary.
@@
"NuHepMC.Beam[PDG].MonoEnergetic.Energy" can be used to signal the beam energy in the lab frame, but the usage of this attribute is optional as the energy can be determined from the first (or any) event in the vector.For the case of a "Histogram"-type distribution, the histogram should be encoded into two HepMC3::VectorDoubleAttribute per beam species on the HepMC3::GenRunInfo:
@@
should be used to signal that the number of neutrinos in a given histogram is found by multiplying the bin content by the bin width, rather than from the content alone. While this might be determined by parsing the RateUnit attribute, existing neutrino generators make different assumptions when sampling input neutrino beam energy distributions, so we specify an explicit attribute. If this attribute is not provided, then it is expected that the number of neutrinos in a given bin is specified by the bin content alone and is independent of the width of the bin.
For a suggestion on how to encode useful information about more realistic neutrino beam descriptions, see E.S.1.
G.C.8
@@
"NuHepMC.AdditionalParticleNumbers" attribute, according to G.R.8, the HepMC3::GenRunInfo instance may contain an attribute giving a description of the particle:-
@@
For non-standard particles that should be further simulated by particle propagation simulations, such as GEANT4 [27], additional information encoded here may be enough to enable automatic propagation. In this version of NuHepMC, we do not attempt to prescribe a format for such information but highlight that HepMC3::GenRunInfo attributes of the form,
"NuHepMC.AdditionalParticleNumber[PDG].SimName.AttrName", might be a useful for communicating such additional information. These additional attributes should include, at a minimum, the particle’s mass, width, spin, and electric charge. -

### v3

@@ 1 Introduction
Recent discussions about the future of neutrino research, both at workshops focused on simulations of neutrino interactions [3] and as part of the HEP-wide Snowmass 2021 community planning process [4], have emphasized the need for greater flexibility in the use of neutrino event generators and related software. A particularly problematic technical barrier is the current lack of a common data format for representing the output events: lists of simulated particles involved in a neutrino interaction together with information describing their properties and relationships. At present, each neutrino event generator group maintains a unique output format, substantially complicating the use of multiple generators in large-scale experimental simulation workflows [5, 6]. The adoption of standard formats in the collider community has streamlined many components of the analysis pipeline. This has enabled straightforward analysis preservation in tools like Rivet [7, 8, 9], allowed theorists to reinterpret experimental limits on one exotic physics scenario in light of others [10, 11], simplified comparisons between generators (see, e.g., Ref. [12]), and supported interoperability between simulation tools [13, 14, 15].
@@
In Sec. 2, we present the specification of the NuHepMC format.111For the most up-to-date specification see https://github.com/NuHepMC/Spec. The flux-averaged total cross section, a particularly important quantity for analyzing neutrino scattering simulations, is discussed in Sec. 3. Finally, Sec. 4 uses NuHepMC as an output format for GENIE, NuWro, NEUT, ACHILLES [24], and MARLEY [25] as well as an input format in NUISANCE to demonstrate a first NuHepMC-based analysis of neutrino event generator predictions.
@@ 2 Specification
The details of the NuHepMC specification are broken down into three categories that describe four components from HepMC3. Each element of the specification is labeled as a Requirement, a Convention, or a Suggestion. The requirements dictate a minimum level of information to be included when writing out events. The conventions are optional details that an event generator group can choose to include or omit while still conforming to the NuHepMC standard.222Helper functions and utilities for writing, reading, and interpreting NuHepMC events in C ++ or Python can be found at https://github.com/NuHepMC/cpputils. Finally, the suggestions are optional recommendations that are less strongly encouraged than the conventions.
@@ 2.1 Labeling scheme
The elements of the specification are enumerated below in the form
<Component>.<Category>.<Index>, where the component of interest is given as G for generator
run metadata, E for event metadata, V for vertex information, and P for
particle information. The category is denoted by R for a requirement, C for a
convention, and S for a suggestion. For example, the second convention for
event metadata would be labeled as E.C.2.
If conventions or suggestions prove useful, become widely adopted, and are considered stable, they may become requirements in future versions of this specification. G.R.4 defines a requirement for event generator authors to signal whether or not specific optional elements of this specification have been followed.
@@ 2.2 HepMC3 C++ classes
Throughout this standard, references are made to various HepMC3 C
classes,
e.g.,
HepMC3::GenRunInfo. However, these are used as a convenient
handle for data objects. This specification should not be considered specific
to the HepMC3 C
reference implementation.
In principle, the HepMC3 ASCII format can be written out to file without any required dependencies. However, it is important to note that the HepMC3 library does provide bindings for both C ++ and Python. The implementation within each tool using the NuHepMC format is left to the developers. At the time of writing there are official or unofficial conversion tools available for all of the aforementioned neutrino event generators, see Section 4.
@@ 2.3.1 Requirements
All NuHepMC vectors333The term vector is used herein, rather than file, as HepMC3 events are frequently piped or streamed between MC tools without ever being wholly persisted to memory or disk. must contain a HepMC3::GenRunInfo instance. -
@@
A NuHepMC HepMC3::GenRunInfo instance must contain the following three
HepMC3::IntAttributes that specify the version of NuHepMC that is implemented:-
"NuHepMC.Version.Major"
"NuHepMC.Version.Minor"
"NuHepMC.Version.Patch"
This document describes version 1.0.0 of NuHepMC.444A script is provided for converting existing NuHepMC files from an older version of the standard to the latest version of the standard at https://github.com/NuHepMC/Spec/scripts/NuHepMC_update_standard.py
@@ A NuHepMC HepMC3::GenRunInfo instance must contain a
Signaling Followed Conventions:
To signal to a user that an implementation follows a named convention from this specification, a HepMC3::VectorStringAttribute attribute must be added to the HepMC3::GenRunInfo instance named "NuHepMC.Conventions" containing the labels of the conventions adhered to, e.g. "G.C.2". n.b. that G.R.5 requires the implementation of at least one convention and so this attribute must exist. -
G.R.5
Flux-Averaged Total Cross Section:
The flux-averaged total cross section, , is a scaling factor that is needed to convert a distribution of simulated events into a cross-section prediction. Details on the mathematical definition of this quantity are given in Sec. 3.1. If the value of is known at the start of a generator run, convention G.C.2 should be used to store it in the HepMC3::GenRunInfo.In general, the value of is not known at the start of a generator run, but it can be calculated as events are produced (see Appendix A). In this case, convention E.C.4 should be used to store the running estimate and associated statistical uncertainty in each HepMC3::GenEvent.
One of these two methods for communicating to users must be implemented and signaled via G.R.4.
G.R.6
Cross Section Units and Target Scaling:
There are a variety of units typically used to report both measured and predicted cross sections in HEP. For neutrino cross sections specifically, per nucleon is common, but not ubiquitous. Both of the following
HepMC3::StringAttributes must be included on the HepMC3::GenRunInfo to indicate the cross section units used within a vector. Possible values of the attributes are not restricted by this specification, but the meanings of the following values are reserved to standardize existing conventions. It is strongly recommended that new implementations use these wherever possible.-
"NuHepMC.Units.CrossSection.Unit":
"pb": Picobarns or . Our recommendation.
"1e-38 cm2": The choice of is the most frequent in the neutrino literature.
"NuHepMC.Units.CrossSection.TargetScale":
"PerAtom": Our recommendation. For this choice, the target number density mentioned in Sec. 3 is expressed in atoms per unit volume.
"PerNucleon": A common alternative in the literature.
The chosen units should be assumed to apply to the reported value of , which is a property of the generator run and may be stored according to either G.C.2 or E.C.4. Consistent units should also be used for the total (E.C.2) and process-specific (E.C.3) cross sections describing the primary interaction in each individual event.
It is ultimately up to the user to parse these attributes and decide whether any additional scaling is needed for their purposes, but the use of the above reserved values will facilitate automated processing. The recommended value of picobarns per atom is chosen to remain consistent with the assumptions of other tools that read and write HepMC3 files beyond the neutrino community, such as Rivet [7, 8, 9]. Consistency facilitates interoperability and reduces maintenance burdens.
G.R.7
Event Weights:
For weights that will be calculated for every event, HepMC3 provides an interface for storing the weight names only once in the HepMC3::GenRunInfo instance. At least one event weight named "CV" must be declared in the
HepMC3::GenRunInfo instance and filled for every event.This weight may be or constant for every event in a generator run (in the case of an unweighted event vector). This weight must always be included by a user when producing cross section predictions from a NuHepMC vector and should never be assumed to be for every event.
G.R.8
@@
A NuHepMC HepMC3::GenRunInfo instance must contain a
HepMC3::VectorIntAttribute named "NuHepMC.ProcessIDs" listing all physics process IDs as integers. For each valid process ID, the
HepMC3::GenRunInfo instance must also contain two other attributes giving a name and description of each:-
@@
where ID enumerates all process IDs present in "NuHepMC.ProcessIDs". (See also E.R.3).
G.R.9
@@
HepMC3::VectorIntAttribute named
"NuHepMC.VertexStatusIDs" declaring any generator-specific status codes used. Including the standard HepMC3 codes in this list is optional, but they must not be reused to mean something different than in the HepMC3 specification. For each declared vertex status, the HepMC3::GenRunInfo instance must also contain two other attributes giving a name and description of each:-
@@
where ID enumerates all status codes present in
"NuHepMC.VertexStatusIDs" (See also V.R.1). -
G.R.10
@@
HepMC3::VectorIntAttribute named
"NuHepMC.ParticleStatusIDs" declaring any generator-specific status codes used. Including the standard HepMC3 codes in this list is optional, but they must not be reused to mean something different than in the HepMC3 specification. For each valid particle status, the HepMC3::GenRunInfo instance must also contain two other attributes giving a name and description of each:-
@@
where ID enumerates all status codes present in
"NuHepMC.ParticleStatusIDs" (see P.R.1 for more details). -
G.R.11
@@
Essentially all event generators in HEP use a standard set of integer codes to identify particle species. This numbering scheme is maintained by the Particle Data Group (PDG) and is regularly updated in their Review of Particle Physics [27, Sec. 45, p. 733].It is expected that neutrino event generators may need to use codes for non-standard particle species (i.e., those without an existing PDG code) for a variety of applications. This could include simulating exotic physics processes involving new particles as well as implementing bookkeeping methods involving generator-specific pseudoparticles.
@@
HepMC3::VectorIntAttribute named
"NuHepMC.AdditionalParticleNumbers" declaring any particle codes used that are not defined in the current PDG numbering scheme. Including any of the standard codes in this list is permitted but not required. The standard particle codes must not be reused to mean something different than in the PDG specification.For each additional particle code, the HepMC3::GenRunInfo instance must also contain an attribute giving a unique name to the represented particle species:
@@
"NuHepMC.AdditionalParticleNumbers".See also G.C.5 for a suggested way of storing descriptions of these special particle species.
2.3.2 Conventions
@@
Vector Exposure:
Each vector should contain a description of the exposure of the generator run. When simulating with some experimental exposure, often represented for accelerator neutrino experiments in units of “protons on target” (POT), the exposure should be described. Two attributes are reserved for signaling the exposure to users. One or both can be provided.-
type: HepMC3::DoubleAttribute, name: "NuHepMC.Exposure.POT"
type: HepMC3::DoubleAttribute,
name: "NuHepMC.Exposure.Livetime"
@@
Flux-averaged Total Cross Section:
If known at the start of a run, the value of should be stored as a
HepMC3::DoubleAttribute in the HepMC3::GenRunInfo named
"NuHepMC.FluxAveragedTotalCrossSection". Optionally, the uncertainty in the flux-averaged total cross section may be stored as a
HepMC3::DoubleAttribute in the HepMC3::GenRunInfo named
"NuHepMC.FluxAveragedTotalCrossSectionUncertainty". -
@@
Modeling components implemented based on published work should always be fully cited. The HepMC3::GenRunInfo should contain at least one
HepMC3::VectorStringAttribute for each relevant modeling component, named according to the pattern "NuHepMC.Citations.Comp.Type". Valid substitutions for the Comp and Type fields are not restricted by this standard beyond the requirement that they are pure mixed-case alpha-numeric. The keys Comp=Generator and Comp=Process[ID] are reserved for use in specifying the main citation(s) for the interaction generator and for citing theoretical motivations behind individual processes, respectively. For common reference formats in the HEP field, some reserved values for the Type field are:-
@@
A tool that can read this metadata and automatically produce a BibTeX file containing entries for all cited publications is briefly introduced in Appendix C.
G.C.4
@@
Each vector should contain a description of the incident flux of beam particles used to simulate the events. For many MC studies and experimental simulations in which the detector is not physically close to the source, a simple univariate energy distribution is enough to describe the particle beam. The two types of energy distribution covered by this convention are mono-energetic beams and those with energy spectra that can be described by one-dimensional histograms. The type should be signaled via a HepMC3::StringAttribute named "NuHepMC.Beam.Type" with value
"MonoEnergetic" or "Histogram" stored on the HepMC3::GenRunInfo. For both types, relevant units can be signaled via two attributes:-
"NuHepMC.Beam.EnergyUnit". Possible values of the attribute are not restricted, but the meanings of "MEV" and "GEV" are reserved. This attribute should always exist and be not empty.
"NuHepMC.Beam.FluxUnit". Possible values of the attribute are not restricted, but the meaning of "Arbitrary" is reserved to signal that the normalization of the distribution was not known or used by the simulation. If this attribute is not used then the normalization will be assumed to be arbitrary.
@@
"NuHepMC.Beam[PDG].MonoEnergetic.Energy" can be used to signal the beam energy in the lab frame, but the usage of this attribute is optional as the energy can be determined from the first (or any) event in the vector.For the case of a "Histogram"-type distribution, the histogram should be encoded into two HepMC3::VectorDoubleAttribute per beam species on the
HepMC3::GenRunInfo:-
@@
should be set to true to indicate that the total neutrino flux in a bin of the histogram should be computed by multiplying the bin content by the bin width. A value of false, assumed by default when this attribute is not present, signals that the bin content should be used alone. While this distinction might be determined by carefully parsing the FluxUnit attribute, existing neutrino generators make different assumptions when sampling from input neutrino beam energy distributions. An explicit attribute is therefore defined here to reduce ambiguity.
For a suggestion on how to encode additional information about more realistic neutrino beam descriptions, see E.S.1.
General flux handling is a complex problem that is not currently standardized. This convention is suitable for simplified studies by theorists and phenomenologists interested in accelerator-based neutrino experiments. A standardized flux format usable for many other applications, e.g., atmospheric neutrino measurements, as well as its corresponding description within the event record, is outside the scope of this specification. Development of such a standard is left to future community efforts.
G.C.5
@@
"NuHepMC.AdditionalParticleNumbers" attribute, according to G.R.11, the HepMC3::GenRunInfo instance may contain an attribute giving a description of the particle:-
@@
For non-standard particles that should be further simulated by particle propagation simulations, such as GEANT4 [28], additional information encoded here may be enough to enable automatic propagation. In this version of NuHepMC, no attempt is made to standardize a format for such information, but it is suggested that HepMC3::GenRunInfo attributes of the form,
"NuHepMC.AdditionalParticleNumber[PDG].SimName.AttrName", may be useful and could be standardized in future versions of this specification. These additional attributes could be expected to include, at minimum, the particle’s mass, width, spin, and electric charge. -
2.3.3 Suggestions